OpenStudy (anonymous):

A spherical party balloon is being inflated with helium pumped in at a rate of 4 cubic feet per minute. How fast is the radius growing at the instant when the radius has reached 2 ft? The volume of a sphere of radius r is V =4/3(3.14)r^3

OpenStudy (anonymous):

Where are you in relationship to earths core?

OpenStudy (anonymous):

does that even matter?

OpenStudy (anonymous):

Yes, good question^^.

OpenStudy (anonymous):

this is RELATED RATES, calculus

OpenStudy (anonymous):

Yes, it does. Without it your lacking the proper centrifugal force.

OpenStudy (anonymous):

can someone help me with this question

OpenStudy (anonymous):

I just need to do this for practice. Hold on, let me try it on the whiteboard.

OpenStudy (anonymous):

dv/dt=4/3*pi*r^2*(dr/dt) dv/dt=4, plug that in now solve for dr/dt

OpenStudy (anonymous):

and plug in r = 2

OpenStudy (anonymous):

i thought u already plugged in 2

OpenStudy (anonymous):

sorry dv/dt = 4*pi*r^2*dr/dt

OpenStudy (anonymous):

messed up derivative, my bad

OpenStudy (anonymous):

you have to plug in the value for r after taking the derivative to that the rate you get (dr/dt) is solved for, right when r = 2

OpenStudy (anonymous):

how would i find the derivative of \[4Pir^2\]

OpenStudy (anonymous):

16(pi)r?

OpenStudy (anonymous):

v=(4/3)*pi*r^3 the only things changing with respect to time in this equation are the radius and the volume. so when you derive this equation you will derive both of these with respect to time. (4/3)*pi are just constants so d/dt of r^3 = 3*r^2 which leads to dv/dt=4*pi*r^2*dr/dt

OpenStudy (anonymous):

sorry d/dt of r^3 = 3*r^2* (dr/dt),

OpenStudy (anonymous):

then find the derivative of that ^^^

OpenStudy (anonymous):

6r?

OpenStudy (anonymous):